S R I: Interface Box

If we connect a 0.5 F capacitor in parallel with the capacitor shown in Fig. 4.1, the time constant of the circuit

A. No longer depends on the value of R_ee

B. Increases

C. Decreases

D. Does not change
Answer #1

Time Constant of RC network shown is

T=Req *C

now if another capacitor 0.5 fard is in parallel with first capacitor net capacitance =(C+0.5)

So T=Req *(C+0.5)

THUS TIME CONSTANT INCRESAES (B)